# Supplemetary Material for the paper

### v1

@@ 7 Conclusion
We introduce a novel overparameterized factorization based on the Hadamard product to induce IR for structured sparse recovery of MMV. We show that the optimization dynamics naturally induce an implicit bias toward row-sparse solutions. Our characterization of the row norm dynamics identifies a momentum-like effect, where rows with larger norms grow significantly faster than those with smaller norms, leading to an incremental learning process that provably favors low-rank solutions. The theoretical analysis, validated through simulations, establishes two key properties of the optimization dynamics. We prove that the relative scaling between the parameters is preserved both globally and for each row throughout the optimization process. We also provide a formal convergence guarantee that with a sufficiently small and balanced initialization, the algorithm’s trajectory is guaranteed to remain close to an idealized row-sparse solution. We show that the performance of the proposed method is comparable to the traditional algorithms like M-OMP, M-SP and M-FOCUSS. Critically, the performance of the proposed framework is achieved without requiring prior knowledge of the sparsity level of the signal or its noise variance. This research successfully bridges the gap between the theoretical understanding of implicit regularization and its practical application, offering a versatile and tuning-free tool for the complex domain of MMV recovery.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] See Definition 3.1 and Theorem 5.1 statement.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] See section 5
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes] See appendix B and C
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] See section E
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] See section 6
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] See section D
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Appendix A Gradient Updates
This section details the derivation of the gradient descent update rules for the parameters $\mathbf{g}$ and $\mathbf{V}$. We consider the loss function for the proposed parameterization, $\mathbf{X}=\mathbf{g}^{\odot 2}\mathbf{1}_{L}\odot\mathbf{V}$, which is reproduced here for readability:
$\mathcal{L}(\mathbf{g},\mathbf{V})\ =\|\mathbf{Y}-\mathbf{A}(\mathbf{g}^{\odot 2}\mathbf{1}_{L}\odot\mathbf{V})\|_{F}^{2}.$ (20)
Let $\boldsymbol{\Lambda}:=\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}\mathbf{X}\right)$. We begin by deriving the partial derivative of the loss function $\mathcal{L}(\mathbf{g},\mathbf{V})$ with respect to a single element $g_{l}$ (the $l$-th element of $\mathbf{g}$):
@@
$=-4\sum\limits_{i\in[M]}\sum\limits_{j\in[L]}\left(Y_{ij}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}V_{kj}\right)A_{il}g_{l}V_{lj}$ (22)
$=-4g_{l}\sum\limits_{i\in[M]}\sum\limits_{j\in[L]}A_{il}\left(Y_{ij}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}V_{kj}\right)V_{lj}$ (23)
$=-4g_{l}\sum\limits_{j\in[L]}\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}(\mathbf{g}^{\odot 2}\mathbf{1}_{L}\mathbf{V})\right)_{lj}V_{lj}$ (24)
$=-4g_{l}\sum\limits_{j\in[L]}\Lambda_{lj}V_{lj}.$ (25)
@@
$\nabla_{\mathbf{g}}\mathcal{L}(\mathbf{g},\mathbf{V})=-4\mathbf{g}\odot\left((\boldsymbol{\Lambda}\odot\mathbf{V})\mathbf{1}_{L}^{\top}\right).$ (26)
Similarly, we compute the gradient of the loss function with respect to the element $V_{lm}$ (the $(l,m)$-th element of $\mathbf{V})$:
$\nabla_{V_{lm}}$ $\mathcal{L}(\mathbf{g},\mathbf{V})=\nabla_{V_{lm}}\sum\limits_{i\in[M]}\sum\limits_{j}\left(Y_{ij}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}V_{kj}\right)^{2}$ (27)
$=-2\sum\limits_{i\in[M]}\left(Y_{im}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}V_{km}\right)A_{il}g_{l}^{2}$ (28)
$=-2g_{l}^{2}\sum\limits_{i\in[M]}A_{il}\left(Y_{im}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}V_{km}\right)$ (29)
$=-2g_{l}^{2}\left(\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}(\mathbf{g}^{\odot 2}\mathbf{1}_{L}\odot\mathbf{V}\right)\right)_{lm}$ (30)
$=-2g_{l}^{2}\Lambda_{lm}.$ (31)
@@
$\nabla_{\mathbf{V}}\mathcal{L}(\mathbf{g},\mathbf{V})=-2\mathbf{g}^{\odot 2}\mathbf{1}_{L}\odot\boldsymbol{\Lambda}$ (32)
@@ Appendix B Characterization of the gradient flow
For the system evolving under continuous gradient flow for the set of updates (4), (5) and (6), derived from the loss function $\mathcal{L}(\mathbf{g},\mathbf{V})=\|\mathbf{Y}-\mathbf{A}(\mathbf{g}^{\odot 2}\mathbf{1}_{L}\mathbf{V})\|_{F}^{2}$, the following properties hold for any time $t\geq 0$:
@@
    Global balancedness: The quantity $\frac{1}{2}\left\|\mathbf{g}(t)_{2}^{2}\right\|-\left\|\mathbf{V}(t)_{F}^{2}\right\|$ is conserved throughout the optimization process:
@@
Lemma B.3 (Lemma 5.2 restated).
@@
if $\left<\boldsymbol{\lambda}_{i}(t),\hat{\mathbf{x}}_{i}(t)\right><0$:
@@
Lemma B.4 (Lemma 5.3 restated).
@@
For the system evolving under continuous gradient flow, as described in Lemma B.5:
@@ 2. Case II:
This system forms an initial value problem in both $g_{i}(t)$ and $V_{il}(t)$. Because the loss function gradient is locally Lipschitz continuous with respect to the parameters $(\mathbf{g}(t),\mathbf{V}(t))$ and uniformly continuous with respect to time $t$, Lemma B.5 and teschl2012ordinary111(teschl2012ordinary, Theorem 2.2) (Picard-Lindelöf): Suppose $f\in C\left(U,\mathbb{R}^{n}\right)$, where $U$ is an open subset of $\mathbb{R}^{n+1}$, and $\left(t_{0},x_{0}\right)\in U$. If $f$ is locally Lipschitz continuous in the second argument, uniformly with respect to the first, then there exists a unique local solution $\bar{x}(t)\in C^{1}(I)$ of the initial value problem, where I is some interval around $t_{0}$., guarantees a unique solution for $g_{i}(t)$ and $V_{il}(t)$. At $t=0$,
@@
As all the necessary conditions of razin2021implicit are met, we have $g_{i}(t)>0$, $\|V_{i:}(t)\|>0$. ∎
@@
Appendix C Supporting Lemmas and the Proof of the Main Theorem
@@
$\tilde{g}_{i}(0)=\begin{cases}\sqrt{2}\rho^{1/3}\ &\forall i\in\mathcal{S}_{K}\\ 0&\forall i\notin\mathcal{S}_{K}\end{cases}\quad;\;\;\quad\tilde{V}_{i,j}(0)=\begin{cases}\rho^{1/3}\frac{V_{ij}(0)}{\|\mathbf{V}_{i:}(0)\|}\ &\forall i\in\mathcal{S}_{K}\\ 0&\forall i\notin\mathcal{S}_{K}\end{cases},\ \forall\ j\in[N],$ (114)
which results in
$\|\tilde{\mathbf{V}}_{i:}(0)\|_{2}=\begin{cases}\rho^{1/3}\ &\forall i\in\mathcal{S}_{K},\\ 0&\forall i\notin\mathcal{S}_{K}\end{cases}$ (115)

### v2

@@ 7 Conclusion
We introduce a novel overparameterized factorization based on the Hadamard product to induce IR for structured sparse recovery of MMV. We show that the optimization dynamics naturally induce an implicit bias toward row-sparse solutions. Our characterization of the row norm dynamics identifies a momentum-like effect, where rows with larger norms grow significantly faster than those with smaller norms, leading to an incremental learning process that provably favors low-rank solutions. We prove that the relative scaling between the parameters is preserved both globally and for each row throughout the optimization process. Furthermore, by employing a Lyapunov-based analysis of the gradient flow, we derive a formal convergence guarantee showing that with a sufficiently small and balanced initialization, the algorithm’s trajectory remains close to an idealized row-sparse solution. We demonstrate that the proposed method achieves performance comparable to algorithms like M-OMP, M-SP, AMP-MMV and M-FOCUSS when they are provided with exact parameters. Moreover, our framework significantly outperforms these baselines in scenarios where they suffer from parameter mismatch, and maintains robust recovery even in low SNR regimes where M-BSBL performs poorly. Critically, the performance of the proposed framework is achieved without requiring prior knowledge of the sparsity level of the signal or its noise variance.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Gradient Updates
This section details the derivation of the gradient descent update rules for the parameters $\mathbf{g}(t)$ and $\mathbf{V}(t)$. We consider the loss function for the proposed parameterization, $\mathbf{X}(t)=\mathbf{g}^{\odot 2}(t)\mathbf{1}_{L}\odot\mathbf{V}(t)$, which is reproduced here for readability:
$\mathcal{L}(\mathbf{g}(t),\mathbf{V}(t))\ =\|\mathbf{Y}-\mathbf{A}(\mathbf{g}^{\odot 2}(t)\mathbf{1}_{L}\odot\mathbf{V}(t))\|_{F}^{2}.$ (20)
Let $\boldsymbol{\Lambda}:=\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}\mathbf{X}(t)\right)$. We begin by deriving the partial derivative of the loss function $\mathcal{L}(\mathbf{g}(t),\mathbf{V}(t))$ with respect to a single element $g_{l}(t)$ (the $l$-th element of $\mathbf{g}(t)$):
@@
$=-4\sum\limits_{i\in[M]}\sum\limits_{j\in[L]}\left(Y_{ij}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}(t)V_{kj}(t)\right)A_{il}g_{l}(t)V_{lj}(t)$ (22)
$=-4g_{l}(t)\sum\limits_{i\in[M]}\sum\limits_{j\in[L]}A_{il}\left(Y_{ij}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}(t)V_{kj}(t)\right)V_{lj}(t)$ (23)
$=-4g_{l}(t)\sum\limits_{j\in[L]}\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}(\mathbf{g}^{\odot 2}(t)\mathbf{1}_{L}\mathbf{V}(t))\right)_{lj}V_{lj}(t)$ (24)
$=-4g_{l}(t)\sum\limits_{j\in[L]}\Lambda_{lj}(t)V_{lj}(t).$ (25)
@@
$\nabla_{\mathbf{g}(t)}\mathcal{L}(\mathbf{g}(t),\mathbf{V}(t))=-4\mathbf{g}(t)\odot\left((\boldsymbol{\Lambda}(t)\odot\mathbf{V}(t))\mathbf{1}_{L}^{\top}\right).$ (26)
Similarly, we compute the gradient of the loss function with respect to the element $V_{lm}(t)$ (the $(l,m)$-th element of $\mathbf{V}(t))$:
$\nabla_{V_{lm}(t)}$ $\mathcal{L}(\mathbf{g}(t),\mathbf{V}(t))=\nabla_{V_{lm}(t)}\sum\limits_{i\in[M]}\sum\limits_{j}\left(Y_{ij}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}(t)V_{kj}(t)\right)^{2}$ (27)
$=-2\sum\limits_{i\in[M]}\left(Y_{im}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}(t)V_{km}(t)\right)A_{il}g_{l}^{2}(t)$ (28)
$=-2g_{l}^{2}(t)\sum\limits_{i\in[M]}A_{il}\left(Y_{im}-\sum\limits_{k\in[N]}A_{ik}g_{k}^{2}(t)V_{km}(t)\right)$ (29)
$=-2g_{l}^{2}(t)\left(\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}(\mathbf{g}^{\odot 2}(t)\mathbf{1}_{L}\odot\mathbf{V}(t)\right)\right)_{lm}$ (30)
$=-2g_{l}^{2}(t)\Lambda_{lm}(t).$ (31)
@@
$\nabla_{\mathbf{V}(t)}\mathcal{L}(\mathbf{g}(t),\mathbf{V}(t))=-2\mathbf{g}^{\odot 2}(t)\mathbf{1}_{L}\odot\boldsymbol{\Lambda}(t)$ (32)
@@ Appendix B Characterization of the gradient flow
For the system evolving under continuous gradient flow for the set of update equations (4), (5) and (6), derived from the loss function $\mathcal{L}(\mathbf{g}(t),\mathbf{V}(t))=\|\mathbf{Y}-\mathbf{A}(\mathbf{g}(t)^{\odot 2}\mathbf{1}_{L}\odot\mathbf{V}(t))\|_{F}^{2}$, the following properties hold for any time $t\geq 0$:
@@
    Global balancedness: The quantity $\frac{1}{2}\left\|\mathbf{g}(t)\right\|_{2}^{2}-\left\|\mathbf{V}(t)\right\|_{F}^{2}$ is conserved throughout the optimization process:
@@
Theorem B.4 (Theorem 5.3 restated).
@@
if $\left<\boldsymbol{\lambda}_{i}(t),\hat{\mathbf{x}}_{i}(t)\right><0$,
@@
Corollary B.5 (Corollary 5.4 restated).
@@
For the system evolving under continuous gradient flow, as described in Corollary B.5:
@@ 2. Case II:
This system forms an initial value problem in both $g_{i}(t)$ and $V_{il}(t)$. Because the loss function gradient is locally Lipschitz continuous with respect to the parameters $(\mathbf{g}(t),\mathbf{V}(t))$ and uniformly continuous with respect to time $t$, Lemma B.6 and Teschl (2012, Theorem 2.2)111(Teschl, 2012, Theorem 2.2) (Picard-Lindelöf): Suppose $f\in C\left(U,\mathbb{R}^{n}\right)$, where $U$ is an open subset of $\mathbb{R}^{n+1}$, and $\left(t_{0},x_{0}\right)\in U$. If $f$ is locally Lipschitz continuous in the second argument, uniformly with respect to the first, then there exists a unique local solution $\bar{x}(t)\in C^{1}(I)$ of the initial value problem, where I is some interval around $t_{0}$., guarantees a unique solution for $g_{i}(t)$ and $V_{il}(t)$. At $t=0$,
@@
As all the necessary conditions of Razin et al. (2021, Lemma 5) are met, we have $g_{i}(t)>0$, $\|V_{i:}(t)\|>0$. ∎
@@
Appendix C Supporting Lemmas and the Proof of the Theorem 5.6
@@
$\tilde{g}_{i}(0)=\begin{cases}\sqrt{2}\rho^{1/3}\ &\forall i\in\mathcal{S}_{K}\\ 0&\forall i\notin\mathcal{S}_{K}\end{cases}\quad;\;\;\quad\tilde{V}_{i,j}(0)=\begin{cases}\frac{1}{\sqrt{L}}\rho^{1/3}\ &\forall i\in\mathcal{S}_{K}\\ 0&\forall i\notin\mathcal{S}_{K}\end{cases},\ \forall\ j\in[N].$ (114)
Consequently, the initial row norms are given by
$\|\tilde{\mathbf{V}}_{i:}(0)\|_{2}=\begin{cases}\rho^{1/3}\ &\forall i\in\mathcal{S}_{K}\\ 0&\forall i\notin\mathcal{S}_{K}\end{cases}.$ (115)